# Solving the Gross-Pitaevskii Equation with Quantic Tensor Trains: Ground States and Nonlinear Dynamics

### v2

@@ 4.1 BEC in a harmonic potential
We first benchmark our method on a system of a BEC in a harmonic trap. This corresponds to $\Omega=0$ and $g=100$ in the Hamiltonian in Eq. 3. The particle density of the ground-state wavefunction is shown in the inset of Fig. 3(a). This function can be efficiently represented in a QTT format with small bond dimensions. In Fig. 3(a), we show the errors of the QTT approximations with different bond dimensions $D$. The error is defined by $1-|\langle\psi|\psi_{\text{QTT}}\rangle|$ where $|\psi\rangle$ is the exact ground state and $|\psi_{\text{QTT}}\rangle$ is the QTT wavefunction after compression. One can see that QTT achieves machine accuracy with error $\approx 10^{-16}$ with a bond dimension $D=20$, and the error is not sensitive to the number of qubits. We thus choose $D=20$ in the following benchmark.
@@
Figure 3: Errors in the QTT compression of ground-state wavefunctions for (a) a BEC in a harmonic trap and (b) a BEC with vortices, plotted as functions of bond dimension $D$. The inset in panel (a) shows the density profile of the corresponding ground-state wavefunction.
@@
To demonstrate the efficiency of QTT in terms of the number of grid points, we compare the computational time required for a single ITE step between the QTT method and the conventional method. With QTT formats, the number of qubits increases exponentially as the number of qubits increases linearly. This is directly reflected in the computational time: When the number of grid points increases exponentially, the QTT computational time scales only linearly, while the computational time for conventional discretizations grows exponentially, as shown in Fig. 4.
@@
Figure 4: CPU times for a single ITE step using the QTT and conventional methods, plotted as functions of the number of qubits (equivalently, the number of grid points). Results are shown in (a) linear and (b) logarithmic scales to highlight the different scaling behaviors of the two approaches.
@@
We then compare the efficiencies for the two ground-state algorithms, ITE and gradient descent. In Figs. 5(a) and (b), we show the errors of the chemical potential $\mu$ as functions of the CPU time and of the optimization steps for each method. Note that the variational method minimizes the energy $E[\psi]$ rather than the chemical potential $\mu$. Therefore, $\mu$ is not variational and can be lower than the exact value during optimization. For ITE, an optimization step is defined by an evolution of a time step $\delta\tau$, and for the variational method, an optimization is defined by a complete sweep of optimizations. Both methods obtain $\mu$ with an accuracy up to $10^{-10}$ with bond dimension $D=20$, showing excellent representations of the ground state. Although the ITE is slightly faster for an optimization step, gradient descent has a better overall efficiency because, in each iteration, it converges much faster than the ITE. For a fixed $D$, the errors saturate in a large number of optimization steps because the bond dimension $D$ limits the accuracy of $\psi(\mathbf{r})$. One can systematically improve the accuracy by increasing the bond dimension $D$.
@@ 4.2 Rotating BEC
We then consider a more challenging system, a BEC in a rotating frame with angular frequency $\Omega$. The system will induce vortices as the lowest-energy excitations, which form a triangular lattice. The number of vortices depends on the angular momentum $\Omega$ and the system size [47]. Fig. 3(b) shows the errors of QTT compressions for the ground states versus the bond dimension $D$ for different numbers of vortices and with the number of qubits $N_{\mathrm{qub}}=8$ (corresponding to $256$ grid points) for each dimension. Compared to the system without a vortex, the wavefunctions require larger bond dimensions to achieve high accuracies, and the required bond dimensions depend on the number of vortices.
Figure 6: Convergence of the chemical potential $\mu$ computed using gradient descent (GD) and TDVP within the QTT framework, as functions of (a) CPU time and (b) optimization steps, for a BEC ground state containing 7 vortices. Different colors indicate different bond dimensions of the QTT wavefunction. Panel (c) shows the evolution of the bond dimension of $\psi_{\chi}^{2}(\mathbf{r})$ during optimization. The inset displays the corresponding density profile.
@@
Figure 7: Density profiles of the ground states containing (a) 7, (b) 19, (c) 37, and (d) 61 vortices. Panel (e) shows the convergence of the chemical potential for the 19-vortex state, computed using the QTT-based gradient descent (GD) method and a conventional grid-based approach. Dashed lines mark the points at which well-defined triangular vortex lattices begin to emerge.
We then consider several systems with larger $\Omega$, where the numbers of vortices in the ground states range from 19 to 61. We obtain well-converged wavefunctions that show clear triangular lattices in all cases. Their densities are shown in Fig. 7(a)-(d). We found that using the wavefunction with fewer vortices as the initial state significantly improves the convergence when computing ground states with a larger number of vortices. It is also helpful to start from a QTT with fewer qubits and gradually increase the number of qubits by linear interpolation. The corresponding parameters in the Hamiltonian and the bond dimensions of $\psi(\mathbf{r})$ are summarized in Tab. 1.
Number of vortices $\Omega$ $g$ $D$ $N_{\mathrm{qub}}$
7 0.8 100 20 8
19 0.95 100 30 8
37 0.92 500 40 9
61 0.969 500 50 9
Table 1: Hamiltonian parameters $\Omega$ and $g$, the bond dimension $D$, and the number of qubits $N_{\mathrm{qub}}$ for each dimension for the QTT ground states containing different numbers of vortices shown in Fig. 7.
For the system with 19 vortices, we benchmark our gradient descent method with the ITE method with conventional discretization. We compare the errors of chemical potential $\mu$ between the two methods as functions of CPU time, shown in Fig. 7(e). It can be observed that the QTT gradient descent method converges more efficiently than the conventional method. The dashed lines indicate roughly when the well-defined triangular vortex lattices emerge in both methods. The QTT method captures the correct vortex pattern several times faster than the conventional method. We choose this system and parameter set for benchmarking in order to keep the computational time of the conventional method manageable. We expect the QTT method to outperform the conventional method even further in systems with more vortices and finer grids.
@@ 4.3 Dynamics
The real-time evolution of a QTT wavefunction can be performed using the same TDVP algorithm employed for the ground state. Here, we consider the time evolution of a BEC in a harmonic trap ($\Omega=0$) after a sudden quench from an initial interaction strength $g_{1}\approx 12.5$ to a smaller value $g_{2}=g_{1}/2$. After the quench, the system behaves in a so-called breathing mode where the width of the wavefunction oscillates in time with frequency $2\omega$ [58, 47, 59], where $\omega$ is the harmonic trap frequency defined in the Hamiltonian in Eq. 3. Fig. 8(a) shows the wavefunction width as a function of time for two different truncation cutoffs $\epsilon$. In both cases, the results perfectly agree with the expected oscillation frequency.
@@ 5 Conclusion and Discussion
We have introduced an efficient and scalable computational framework based on the QTT format for solving the GPE, which models BECs within the mean-field approximation. By extending tensor network techniques, we adapted the TDVP and developed a gradient-based variational method to handle the nonlinear interaction term. A key element of our approach is the compression of the nonlinear term into a compact QTT representation, which greatly reduces computational complexity without compromising accuracy. This allows for the simulation of large, highly resolved systems that are otherwise difficult to treat with conventional grid-based methods.
@@
Overall, our results establish the QTT framework as a powerful and versatile tool for solving nonlinear quantum equations. By combining high compression efficiency with algorithmic flexibility, it provides a path forward for large-scale simulations in mean-field quantum systems and offers new opportunities for applying tensor network methods to nonlinear and continuum problems in physics.
@@ Acknowledgements
We acknowledge helpful discussions with X. Waintal.
@@
C.-M. C. acknowledges support by MOST (111-2112-M-110-006-MY3 and 113-2112-M-110-026) and the Yushan Young Scholar Program under the Ministry of Education (MOE) in Taiwan. I.-K. Liu is supported the Science and Technology Facilities Council (ST/W001020/1). J.-W. L. acknowledges funding support from the Plan France 2030 ANR-22-PETQ-0007 “EPIQ”
Appendix A Normalization Problem
@@
$\mu_{\mathrm{2D}}=\sum_{i,j}\psi^{*}(\mathbf{r}_{ij})\left[-\frac{1}{2}\nabla_{i,j}^{2}+\frac{1}{2}\omega^{2}\mathbf{r}_{ij}^{2}+g|\psi(\mathbf{r}_{ij})|^{2}-\Omega L_{z}\right]\psi(\mathbf{r}_{ij})\Delta x\Delta y$ (9)
@@
$\sum_{i,j}|\psi(\mathbf{r}_{ij})|^{2}\Delta x\Delta y=1$ (10)
@@
$\tilde{\psi}(\mathbf{r}_{ij})=\psi(\mathbf{r}_{ij})\Delta x$ (11)
@@
$\hat{H}[\tilde{\psi}]=-\frac{1}{2}\nabla^{2}+\frac{1}{2}\omega^{2}r^{2}+\frac{g}{(\Delta x)^{2}}|\tilde{\psi}|^{2}-\Omega L_{z}$ (13)
@@
$\tilde{\mu}_{\mathrm{2D}}=\sum_{i,j}\tilde{\psi}^{*}(\mathbf{r}_{ij})\hat{H}[\tilde{\psi}]\tilde{\psi}(\mathbf{r}_{ij})=\sum_{i,j}\psi^{*}(\mathbf{r}_{ij})\Delta x\left[\hat{H}[\psi]\right]\psi(\mathbf{r}_{ij})\Delta x=\mu_{\mathrm{2D}}$
The rescaled nonlinear coupling parameter exhibits critical scaling behavior:
@@
[…]
@@ Appendix B Review of Quantic Tensor Train
The QTT decomposition represents an $N$-index tensor as a chain of $N$ smaller tensors connected through internal indices known as bond dimensions, as shown in Fig. 9(c). Smooth functions typically have low correlations across different scales, allowing the tensor to be compressed efficiently into a QTT form with small bond dimensions. For certain functions, exact QTT decompositions are known (see Sec. C for examples). In general, however, one must numerically determine the QTT representation using techniques such as tensor cross interpolation (TCI) [60, 61, 62, 63, 64]. TCI is a powerful algorithm that learns QTT tensors from sampled values of the target function, controlling the bond dimensions and thus the accuracy of the approximation.
@@
Similarly, the addition of two functions or operators can be efficiently implemented using standard methods for adding MPSs or MPOs [2]. Using this approach, complex operators can be constructed directly from simpler, elementary operators. For example, the angular momentum operator $\hat{L}_{z}=\hat{x}\hat{p}_{y}-\hat{y}\hat{p}_{x}$ can be built from basic operators $\hat{x}$, $\hat{y}$, $\hat{p}_{x}$, and $\hat{p}_{y}$, each of which has an exact QTTO representation.
@@ Appendix C Exact QTT representations
A linear function $f(x)=ax+b$ can be represented in the QTT form efficiently. It is convenient to first consider $f(x)\to f_{b_{1}b_{2}\cdots b_{N}}=x_{\mathrm{Int}}$ for integer $x_{\mathrm{Int}}$ from 00 to $12^{N}-$ 1. The QTT representation for $f_{b_{1}b_{2}\cdots b_{N}}$ can be easily obtained from the binary representation of integers, $x_{\mathrm{Int}}=\sum_{i=1}^{N}b_{i}2^{i-1}$, where $b_{i}=\left\{0,1\right\}$ are the qubit degrees of freedom and $N$ is the number of qubits. The QTT tensor on site $i$ is thus
@@
– each $\mathbf{t}_{i}$ contributes 00 or $2^{i}$ to $x$ depending on $b_{i}$. The product of the tensors accumulate the sum of $\mathbf{t}_{i}$ to the corner,
@@
Figure 11: The non-zero elements in the tensors $A$, $L$ and $R$ for the operator $\frac{d}{dx}$. The numbers indicate the labels for the non-zero elements. These elements have values equal to one.
@@
$A=\left(\begin{array}[]{cc}\hat{I}&\hat{\sigma}^{+}\\ \hat{0}&\hat{\sigma}^{-}\end{array}\right)$ (20)
@@
Note that in this definition, the derivative of the last point 11$\cdots$1 is defined by the last point 11$\cdots$1 and the first point 00$\cdots$0. In this sense, the wavefunction has periodic boundary condition.

### v3

@@ 4.1 Two-dimensional BEC in a harmonic potential
We first benchmark our method on a system of a BEC in a harmonic trap. This corresponds to $\Omega=0$ and $g=100$ in the Hamiltonian in Eq. 4. The particle density of the ground-state wavefunction is shown in the inset of Fig. 3(a). This function can be efficiently represented in a QTT format with small bond dimensions. In Fig. 3(a), we show the errors of the QTT approximations with different bond dimensions $D$. The fidelity error is defined by $1-|\langle\psi|\psi_{\text{QTT}}\rangle|^{2}$ where $|\psi\rangle$ is the exact ground state and $|\psi_{\text{QTT}}\rangle$ is the QTT wavefunction after compression. One can see that QTT achieves machine accuracy with error $\approx 10^{-16}$ with a bond dimension $D=20$, and the error is not sensitive to the number of qubits. We thus choose $D=20$ in the following benchmark.
@@
Figure 3: Absolute fidelity errors in the QTT of ground-state wavefunctions for (a) a BEC in a harmonic and (b) a BEC with vortices, plotted as functions of bond dimension $D$. The inset in panel (a) shows the density profile of the corresponding ground-state wavefunction. The density profiles of the vortex states can be found in Fig. 7.
@@
To demonstrate the efficiency of QTT in terms of the number of grid points, we compare the computational time required for a single time-evolution step between the QTT method and the explicit ITE method with regular finite-difference discretization. In QTT formats, the number of grid points increases exponentially with the number of qubits, whereas the number of qubits increases linearly. This is directly reflected in the computational time: When the number of grid points increases exponentially, the QTT computational time scales only linearly, while the computational time for conventional discretizations grows exponentially, as shown in Fig. 4.
@@
Figure 4: CPU times for a single ITE step using the QTT and conventional finite-difference methods, as functions of the number of qubits (equivalently, the number of grid points). Results are shown in (a) linear and (b) logarithmic scales to highlight the different scaling behaviors of the two approaches.
@@
We then compare the efficiencies for the two ground-state algorithms, ITE and gradient descent. In Figs. 5(a) and (b), we show the errors of the chemical potential $\mu$ as functions of the CPU time and of the optimization steps for each method. Note that the variational method minimizes the energy $E[\psi]$ rather than the chemical potential $\mu$. Therefore, $\mu$ is not variational and can be lower than the exact value during optimization. For ITE, an optimization step is defined by an evolution of a time step $\delta\tau$, and for the variational method, an optimization is defined by a complete sweep of optimizations. Both methods obtain $\mu$ with an accuracy up to $10^{-10}$ with bond dimension $D=20$, showing excellent representations of the ground state. Although the ITE is slightly faster at each optimization step, gradient descent is overall more efficient because it converges much faster in each iteration. For a fixed $D$, the errors saturate in a large number of optimization steps because the bond dimension $D$ limits the accuracy of $\psi(\mathbf{r})$. One can systematically improve the accuracy by increasing the bond dimension $D$.
@@ 4.2 Rotating BEC
We then consider a more challenging system, a BEC in a rotating frame with angular frequency $\Omega$. The system will induce vortices as the lowest-energy excitations, which form a triangular lattice. The number of vortices depends on the angular momentum $\Omega$ and the system size [pitaevskii2016bose, Cooper01112008], which can be estimated by
$N_{v}\approx\frac{m\Omega}{\pi\hbar}\pi R_{\mathrm{TF}}^{2}$ (10)
where $R_{\mathrm{TF}}=[4gN_{c}/\pi m(\omega^{2}-\Omega^{2})]^{1/4}$ with the condensate number $N_{c}$  [^2] In the dimensionless form in this work, see Appendix A for detail, the dimensionless Thomas-Fermi radius is $R_{\mathrm{TF}}=[4g/\pi(\omega-\Omega^{2})]$ where $g$ is dimensionless and account the contribution of $N_{c}$ according the normalization condition. is the Thomas-Fermi radius of the system. Noting that this estimation ignores the effect of the density inhomogeneity, and hence is only a crude prediction.
Fig. 3(b) shows the errors of QTT compressions for the ground states versus the bond dimension $D$ for different numbers of vortices and with the number of qubits $N_{\mathrm{qub}}=8$ (corresponding to $256$ grid points) for each dimension. Compared to the system without a vortex, the wavefunctions require larger bond dimensions to achieve high accuracy, and the required bond dimensions depend on the number of vortices.
Figure 6: Convergence of the chemical potential $\mu$ computed using gradient descent (GD) and TDVP within the QTT framework, as functions of (a) CPU time and (b) optimization steps, for a BEC ground state containing 7 vortices. Different colors indicate different bond dimensions of the QTT wavefunction. Panel (c) shows the evolution of the bond dimension of $\psi_{\chi}^{2}(\mathbf{r})$ during optimization. The inset displays the corresponding density profile in domain $6,6]^{2}$ with $N_{\mathrm{qub}}=8$.
@@
Figure 7: Density profiles of the ground states containing (a) 7, (b) 19, (c) 37, (d) 61, (e) 91, and (f) 125 vortices. All simulations are performed on a domain of $21,21]^{2}$ using $2^{17}\times 2^{17}$ grid points. The physical parameters, as well as the corresponding energies and chemical potentials of each system, are listed in Tab. 1.
Figure 8: The CPU time of the QTT gradient descent method is compared with several conventional methods, including the backward Euler pseudo-spectral method (BESP), the backward Euler finite-difference method (BEFD), and the explicit ITE method. The system contains 19 vortices, and all simulations start from the same initial state with 7 vortices. The computational domain is $21,21]^{2}$, and the number of grid points is $2^{11}\times 2^{11}$.
To demonstrate the power of the QTT approach, we then consider more challenging systems with larger $\Omega$, where the numbers of vortices in the ground states range from 19 to 125, on a dense $2^{17}\times 2^{17}$ meshgrid. We obtain well-converged wavefunctions that show clear triangular lattices in all cases. Their densities of the wavefunctions are shown in Fig. 7(a)-(f). We find that for large values of $\Omega$, the vortex number can be sensitive to the spatial resolution, making it important to approach the fine-grid limit. In practice, using the converged wavefunction with fewer vortices as the initial state significantly improves convergence when computing ground states with a larger number of vortices. It is also helpful to start from a QTT representation with fewer qubits and gradually increase the number of qubits using linear interpolation. The corresponding Hamiltonian parameters and the bond dimensions of $\psi(\mathbf{r})$ are summarized in Tab. 1.
Number of vortices $\Omega$ $g$ $D$ $N_{\mathrm{qub}}$ $\mu$ $E$
7 0.8 100 30 17 4.35 3.19
19 0.946 100 30 17 2.90 2.25
37 0.913 500 40 17 6.36 4.54
61 0.9616 500 50 17 4.67 3.42
91 0.9617 1100 65 17 6.46 4.59
125 0.972 1550 80 17 6.63 4.68
Table 1: Hamiltonian parameters $\Omega$ and $g$, the bond dimension $D$, and the number of qubits $N_{\mathrm{qub}}$ per dimension for the QTT ground states with different numbers of vortices shown in Fig. 7. The chemical potential $\mu$ and total energy $E$ are evaluated using the gradient descent algorithm.
We benchmark the efficiency of our method on a system with 19 vortices (Fig.7(b)) on a $2^{11}\times 2^{11}$ meshgrids against several conventional methods with regular discretization, including the backward Euler pseudo-spectral method (BESP)[antoine2014gpelab], the backward Euler finite-difference method (BEFD)[antoine2014gpelab], and the explicit ITE method. The initial state is chosen to be a 7-vortex state for all methods. In Fig. 8, we compare the errors in chemical potential $\mu$ and the total energy $E$ as functions of CPU time from different methods. Clearly, the QTT gradient descent method outperforms all the conventional methods. The system parameters are chosen to keep the computational time of the conventional method manageable. The QTT method is expected to be even more efficient than conventional methods on finer grids.
@@ 4.3 Dynamics
The real-time evolution of a QTT wavefunction can be performed using the same TDVP algorithm employed for the ground state. Here, we consider the time evolution of a BEC in a harmonic trap ($\Omega=0$) after a sudden quench from an initial interaction strength $g_{1}\approx 12.5$ to a smaller value $g_{2}=g_{1}/2$. After the quench, the system behaves in a so-called breathing mode where the width of the wavefunction oscillates in time with frequency $2\omega$ [PhysRevB.89.205317, pitaevskii2016bose, zhai2021ultracold], where $\omega$ is the harmonic trap frequency defined in the Hamiltonian in Eq. 4. Fig. 9(a) shows the wavefunction width as a function of time for two different truncation cutoffs $\epsilon$. In both cases, the results perfectly agree with the expected oscillation frequency.
@@ 5 Conclusion and Discussion
We have introduced an efficient and scalable computational framework based on the QTT format for solving the GPE, which models BECs within the mean-field approximation. By extending tensor network techniques, we adapted the ITE and developed a gradient-based variational method to handle the nonlinear interaction term. A key element of our approach is the compression of the nonlinear term into a compact QTT representation, which greatly reduces computational complexity without compromising accuracy. This allows for the simulation of large, highly resolved systems that are otherwise difficult to treat with conventional grid-based methods.
@@
We also note that the QTT framework targets a specific type of approximation and, in principle, can be combined with existing methods, for example, adaptive meshes or implicit time-integration methods. Such combinations would allow approaches that benefit from both techniques. Therefore, QTT should not be viewed simply as competing with other methods, but rather as an alternative approach that can be effectively integrated with them.
Overall, our results establish the QTT framework as a powerful tool for solving nonlinear quantum equations. By combining high compression efficiency with algorithmic flexibility, it provides a path forward for large-scale simulations in mean-field quantum systems and offers new opportunities for applying tensor network methods to nonlinear and continuum problems in physics.
@@ Acknowledgements
We acknowledge helpful discussions with X. Waintal. Chia-Min Chung thanks the CTCP, NYCU, Taiwan, for providing full support.
@@
C.-M. C. acknowledges support by MOST (111-2112-M-110-006-MY3, 113-2112-M-110-026, and 114-2628-M-110-005-MY3) and the Yushan Young Scholar Program under the Ministry of Education (MOE) in Taiwan. I.-K. Liu is supported the Science and Technology Facilities Council (ST/W001020/1). J.-W. L. acknowledges funding support from the Plan France 2030 ANR-22-PETQ-0007 “EPIQ”
Appendix A Normalization
@@
$\mu_{\mathrm{2D}}=\sum_{i,j}\psi^{*}(\mathbf{r}_{ij})\left[-\frac{1}{2}\nabla_{i,j}^{2}+\frac{1}{2}\omega^{2}\mathbf{r}_{ij}^{2}+g|\psi(\mathbf{r}_{ij})|^{2}-\Omega L_{z}\right]\psi(\mathbf{r}_{ij})\Delta x\Delta y,$ (11)
@@
$\sum_{i,j}|\psi(\mathbf{r}_{ij})|^{2}\Delta x\Delta y=1.$ (12)
@@
$\tilde{\psi}(\mathbf{r}_{ij})=\psi(\mathbf{r}_{ij})\Delta x,$ (13)
@@
$\hat{H}[\tilde{\psi}]=-\frac{1}{2}\nabla^{2}+\frac{1}{2}\omega^{2}r^{2}+\frac{g}{(\Delta x)^{2}}|\tilde{\psi}|^{2}-\Omega L_{z}.$ (15)
@@
$\tilde{\mu}_{\mathrm{2D}}=\sum_{i,j}\tilde{\psi}^{*}(\mathbf{r}_{ij})\hat{H}[\tilde{\psi}]\tilde{\psi}(\mathbf{r}_{ij})=\sum_{i,j}\psi^{*}(\mathbf{r}_{ij})\Delta x\left[\hat{H}[\psi]\right]\psi(\mathbf{r}_{ij})\Delta x=\mu_{\mathrm{2D}},$
then, the rescaled nonlinear coupling parameter exhibits critical scaling behavior:
@@
$g^{\prime}>10^{18}\quad\text{(2D case),}\quad g^{\prime}>10^{26}\quad\text{(3D case)}.$ (17)
@@ Appendix B Review of Quantic Tensor Train
The QTT decomposition represents an $N$-index tensor as a chain of $N$ smaller tensors connected through internal indices known as bond dimensions, as shown in Fig. 10(c). Smooth functions typically have low correlations across different scales, allowing the tensor to be compressed efficiently into a QTT form with small bond dimensions. For certain functions, exact QTT decompositions are known (see Appendix C for examples). In general, however, one must numerically determine the QTT representation using techniques such as tensor cross interpolation (TCI) [OSELEDETS201070, 6076873, SAVOSTYANOV2014217, DOLGOV2020106869, fernandez2024learning]. TCI is a powerful algorithm that learns QTT tensors from sampled values of the target function, controlling the bond dimensions and thus the accuracy of the approximation.
@@
Similarly, the addition of two functions or operators can be efficiently implemented using standard methods for adding MPSs or MPOs [SCHOLLWOCK201196]. Using this approach, complex operators can be constructed directly from simpler, elementary operators. For example, the angular momentum operator $\hat{L}_{z}=\hat{x}\hat{p}_{y}-\hat{y}\hat{p}_{x}$ can be built from basic operators $\hat{x}$, $\hat{y}$, $\hat{p}_{x}$, and $\hat{p}_{y}$, each of which has an exact QTTO representation.
@@ Appendix C Exact QTT representations
A linear function $f(x)=ax+b$ can be represented in the QTT form efficiently. It is convenient to first consider $f(x)\to f_{b_{1}b_{2}\cdots b_{N}}=x_{\mathrm{Int}}$ for integer $x_{\mathrm{Int}}$ from 0 to $2^{N}-1$. The QTT representation for $f_{b_{1}b_{2}\cdots b_{N}}$ can be easily obtained from the binary representation of integers, $x_{\mathrm{Int}}=\sum_{i=1}^{N}b_{i}2^{i-1}$, where $b_{i}=\left\{0,1\right\}$ are the qubit degrees of freedom and $N$ is the number of qubits. The QTT tensor on site $i$ is thus
@@
– each $\mathbf{t}_{i}$ contributes 0 or $2^{i}$ to $x$ depending on $b_{i}$. The product of the tensors accumulate the sum of $\mathbf{t}_{i}$ to the corner,
@@
Figure 12: The non-zero elements in the tensors AA, $L$ and $R$ for the operator $\frac{d}{dx}$. The numbers indicate the labels for the non-zero elements. These elements have values equal to one.
@@
$A=\left(\begin{array}[]{cc}\hat{I}&\hat{0}\\ \hat{\sigma}^{+}&\hat{\sigma}^{-}\end{array}\right)$ (22)
@@
Note that in this definition, the derivative of the last point $1\cdots$1 is defined by the last point $1\cdots$1 and the first point $0\cdots$0. In this sense, the wavefunction has periodic boundary condition.